OpenStudy (anonymous):

how do i solve this equation x^4-x^3-x^2-x-2=0 also how do i fine the number and type of root?

OpenStudy (compassionate):

You can solve by adding like terms then factoring it.

OpenStudy (anonymous):

first guess some obvious numbers try \(f(1)\) and if that doesn't work try \(f(-1)\)

OpenStudy (compassionate):

Hmm... @satellite73 , this is a polynomial.

OpenStudy (anonymous):

what grade levle is this????!!!!

OpenStudy (anonymous):

\[f(x)=x^4-x^3-x^2-x-2\] \[f(1)=1-1-1-1-2\neq 0\]

OpenStudy (anonymous):

yes, i know it is a polynomial do you have any idea how to find the zeros of a polynomial of degree 4?

OpenStudy (anonymous):

im a sophomore and its algebra 2

OpenStudy (compassionate):

Of course I do.

OpenStudy (anonymous):

really? love to hear it ok now we know that 1 is not a zero, try \(-1\)

OpenStudy (anonymous):

how would i factor it?

OpenStudy (anonymous):

don't try to factor it until you find the zeros

OpenStudy (anonymous):

\[f(x)=x^4-x^3-x^2-x-2\] \[f(-1)=1+1-1+1-2=0\] got it on the second try!

OpenStudy (anonymous):

now you can factor, since you know \(-1\) is a zero, it must factor as \[x^4-x^3-x^2-x-2=(x+1)(something)\]

OpenStudy (compassionate):

When I did Algebra II, my professor taught us to factor by grouping first. Huh. That's odd.

OpenStudy (anonymous):

find the "something" by division synthetic division is easiest

OpenStudy (anonymous):

we are being taught both synthetic division and factoring. but i don't understand how to group and factor this problem!

OpenStudy (anonymous):

factor by grouping might work if the polynomial has four terms and has been cooked up nicely @madison.hall00 don't try it divide forget factoring

OpenStudy (compassionate):

Yeah. Listen to Satellite. I think the next step is synthetic division and the zero remainders are your roots. Carry on!

OpenStudy (anonymous):

haha thank you guys!